# Harness as an Asset: Enforcing Determinism via the Convergent AI Agent Framework (CAAF)

### 18 Apr 2026 v1

@@ 8.3 Results
System Domain $n$ Correct% Failure Mode Distribution Debate (2 agents, 3 rounds) AD 20 0% Fwd-Viol: 7, Rear-Viol: 13 Debate (2 agents, 3 rounds) Pharma 20 0% Impurity-Viol: 15, Conv-Viol: 4, JSON: 1 Sequential (exec + checker, 3 retries) AD 20 0% Forced: 9, Fwd-Viol: 7, Rear-Viol: 4 Sequential (exec + checker, 3 retries) Pharma 20 0% Impurity-Viol: 7, Conv-Viol: 5, Dual: 2, JSON: 6 CAAF-all-mini (no hint) AD 30 100% Paradox detected: 30/30 CAAF-all-mini (no hint) Pharma 20 100% Paradox detected: 20/20 Mono + UAI (ReAct) AD 20 85% Fwd-Viol: 3 Mono + UAI (ReAct) Pharma 20 95% Impurity-Viol: 1
Finding 9 — Multi-Agent Orchestration Alone is Insufficient. Both the debate and sequential baselines achieve 0% paradox detection across all 80 trials (40 AD + 40 pharma). The debate architecture exhibits a notable failure mode: agents converge on a socially negotiated “consensus” that silently violates one constraint, producing False Consensus—in the pharma domain, 15/20 debate trials converge on impurity-violating solutions where neither agent detects the three-way paradox. The sequential baseline shows even higher noise: 6/20 pharma trials produce unparseable JSON after 3 retries, and the LLM checker approves constraint-violating solutions despite all containing physical violations. Meanwhile, CAAF achieves 100% across both domains. The Mono+UAI ablation achieves 85% on AD and 95% on pharma—substantially better than all baselines, but not perfect, confirming that the UAI assertion engine is the dominant contributor while RAD and State Locking close the remaining gap.
These results confirm that CAAF’s reliability derives from its UAI assertion engine and State Locking mechanism, not from the general principle of multi-agent decomposition. Without a deterministic external validator, adding more agents or more rounds of critique did not bridge the controllability gap at the depths tested (3 debate rounds, 3 checker retries).
Figure 12: Paradox detection rate across architectures and domains. All baselines—monolithic, debate, and sequential—achieve 0% detection in both domains. Only CAAF achieves 100% across both. The Mono+UAI ablation (85% AD, 95% pharma) demonstrates that UAI provides the majority of value, while RAD and State Locking close the remaining gap to 100%.
@@ 9.2 The Economics of Certainty: Redefining Total Cost of Ownership
Table 9: Empirical TCO Comparison
Metric Mono GPT-4o Mono GPT-4o-mini CAAF-all-mini API cost per run (measured) \$0.0145 \$0.0006 \$0.0027 Paradox detection rate 0% 0% 100% Artifact physically verified? No No Yes Estimated HITL cost (per failure) \$50.00 \$50.00 \$5.00 Effective cost per correct artifact $\infty$ $\infty$ \$0.0027
@@
Figure 13: Cost vs. reliability for the no-hint condition (n=30n{=}30). CAAF-all-mini (green) achieves 100% paradox detection at \$0.0027/run, dominating the efficiency frontier. Both monolithic baselines (GPT-4o at \$0.0145/run and GPT-4o-mini at \$0.0006/run) achieve 0% detection, yielding infinite effective cost per correct artifact.
@@ 10.3 Textual Gradients and Optimization
TextGrad (Yuksekgonul et al., 2024) applies the backpropagation metaphor to text, using an LLM evaluator to provide textual gradients for offline prompt and code optimization. CAAF applies a related concept to real-time, zero-shot task execution: the UAI provides a deterministic error trace (analogous to a loss signal), and State Locking ensures that gradient application is monotonically improving. TextGrad’s gradients are fully LLM-generated; CAAF’s gradients are physically anchored by the UAI, making them partially deterministic.
@@ 10.4 Neuro-Symbolic and Formal Methods
Harness Engineering (Böckeler, 2024) established patterns for managing LLM interaction environments to improve reliability in developer productivity tools. CAAF formalizes these patterns into a rigorous framework with guaranteed convergence properties, shifting the application domain from coding assistance (Agent-as-Tool) to safety-critical engineering governance (Agent-as-Controller).
@@ 12 Conclusion
The Mono+UAI ablation (Section 7) reveals that the deterministic UAI assertion engine provides the majority of CAAF’s value, achieving 95% paradox detection even without multi-agent decomposition. The three architectural pillars address complementary failure surfaces: UAI provides deterministic constraint boundaries, RAD ensures input fidelity at scale, and Structured Semantic Gradients with State Locking guarantee monotonic convergence. We hypothesize that RAD’s contribution becomes critical at higher constraint counts where single-pass reasoning degrades—a key axis for future validation.
@@ Appendix C Formal Deadlock Evidence Package
# [SYSTEM DEADLOCK] Formal Paradox Report

### 25 Apr 2026 v2

@@ 8.3 Results
System Domain $n$ Correct% Failure Mode Distribution Debate (2 agents, 3 rounds) AD 20 0% Fwd-Viol: 7, Rear-Viol: 13 Debate (2 agents, 3 rounds) Pharma 20 0% Impurity-Viol: 16, Conv-Viol: 4 Sequential (exec + checker, 3 retries) AD 20 0% Forced: 9, Fwd-Viol: 7, Rear-Viol: 4 Sequential (exec + checker, 3 retries) Pharma 20 0% Impurity-Viol: 9, Conv-Viol: 9, Dual: 2 CAAF-all-mini (no hint) AD 30 100% Paradox detected: 30/30 CAAF-all-mini (no hint) Pharma 20 100% Paradox detected: 20/20 Mono + UAI (prompt-sim) AD 20 85% Fwd-Viol: 3 Mono + UAI (prompt-sim) Pharma 20 95% Impurity-Viol: 1 Mono + UAI (true tool-call, Opus 4 thinking) Pharma 20 100% MUS $\{C1,C2,C4\}$: 20/20 Mono + UAI (true tool-call, Haiku 4.5) Pharma 20 0% Oscillation, max_iters 20/20
Finding 11 — Multi-Agent Orchestration Alone is Insufficient; UAI Is Insufficient Alone at Commodity Cost. Both the debate and sequential baselines achieve 0% paradox detection across all 80 trials (40 AD + 40 Pharma). The debate architecture exhibits a notable failure mode: agents converge on a socially negotiated “consensus” that silently violates one constraint, producing False Consensus—in the Pharma domain, 16/20 debate trials converge on impurity-violating solutions where neither agent detects the three-way paradox. The sequential checker baseline produces a near-even split between impurity and conversion violations (9/20 IMPURITY, 9/20 CONV, 2/20 DUAL), with the LLM checker approving constraint-violating solutions despite all containing physical violations. Meanwhile, CAAF achieves 100% across both domains.
The Mono+UAI ablation yields three disparate data points on Pharma: prompt-sim on GPT-4o-mini reaches 95%; true tool-call on Opus 4 thinking reaches 100%; true tool-call on commodity-tier Haiku 4.5 reaches 0%, with all 20 trials oscillating through max_iters=8 without a paradox declaration (Finding 8). UAI alone therefore closes the controllability gap only given frontier-grade reasoning; at commodity reasoning cost, even faithful tool-call UAI access falls into a stochastic-oscillation trap. The architectural value of CAAF is twofold: enforcement (unlike Mono+UAI, it does not depend on the agent choosing to invoke the assertion engine; unlike multi-agent baselines, it does not depend on consensus dynamics) and oscillation control (the Structured Semantic Gradient with State Locking closes the gap that Haiku+UAI leaves open). See Finding 10 for the cost consequences.
These results confirm that CAAF’s reliability derives from the combination of its UAI assertion engine, Structured Semantic Gradient, and State Locking mechanism, not from the general principle of multi-agent decomposition. Neither UAI grounding alone nor multi-agent orchestration alone bridges the controllability gap; CAAF’s architectural advantage is the composition of these primitives at commodity-model economics.
Figure 12: Paradox detection rate across architectures and domains. All baselines without deterministic UAI—monolithic, debate, and sequential—achieve 0% detection in both domains. Only CAAF achieves 100% across both. Mono+UAI shows large model-dependent variance on Pharma: 95% on GPT-4o-mini (prompt-simulated), 100% on Opus 4 thinking (true tool-call), but 0% on Haiku 4.5 (true tool-call, no thinking) where every run oscillates to max_iters. UAI grounding is necessary but not sufficient at commodity reasoning cost. CAAF closes the oscillation gap structurally via Structured Semantic Gradients with State Locking, reaching 100% on both domains with commodity models.
@@ 9.2 The Economics of Certainty: Redefining Total Cost of Ownership
Table 9: Empirical TCO Comparison. Dashes in the last row indicate the metric is undefined: no correct artifacts were produced, so “cost per correct artifact” is not estimable from the measured data (it would be $\infty$ in the limit of any finite API spend divided by zero successes).
Metric Mono GPT-4o Mono GPT-4o-mini CAAF-all-mini API cost per run (measured) \$0.0145 \$0.0006 \$0.0027 Paradox detection rate 0% 0% 100% Artifact physically verified? No No Yes Estimated HITL cost (per failure) \$50.00 \$50.00 \$5.00 Effective cost per correct artifact —  —  \$0.0027
Not estimable: 0% paradox detection $\Rightarrow$ zero correct artifacts in the denominator.
@@
[…]
9.3 Pharma Cost Comparison at 100% Reliability
[…]
Table 11: Pharma per-trial cost at 100% paradox detection (n=20n{=}20 each, no-hint).
[…]
@@ 10.3 Textual Gradients and Optimization
TextGrad (Yuksekgonul et al., 2024) applies the backpropagation metaphor to text, using an LLM evaluator to provide textual gradients for offline prompt and code optimization. CAAF applies a related concept to real-time, zero-shot task execution: the UAI provides a deterministic error trace (analogous to a loss signal), and State Locking ensures the verified-constraint set is non-regressing across iterations. TextGrad’s gradients are fully LLM-generated; CAAF’s gradients are physically anchored by the UAI, making them partially symbolic-grounded: the Dimension and Direction come from a deterministic boolean failure, while only the Magnitude is LLM-inferred.
@@ 10.4 Neuro-Symbolic and Formal Methods
Harness Engineering. Böckeler (2026) established patterns for managing LLM interaction environments to improve reliability in developer productivity tools. CAAF formalizes these patterns into a rigorous framework with guaranteed convergence properties, shifting the application domain from coding assistance (Agent-as-Tool) to safety-critical engineering governance (Agent-as-Controller).
Concurrent and independent work validates the generality of the Harness Engineering paradigm. Rajasekaran (2026) at Anthropic Labs reports a Generator–Evaluator architecture for long-running application development that is structurally isomorphic to CAAF: a Planner decomposes tasks (cf. RAD), a Generator produces artifacts (cf. Executor), and an Evaluator provides structured feedback (cf. UAI + Reviewer). Their system uses 5–15 iterative cycles with “sprint contracts” that parallel CAAF’s Structured Semantic Gradients, and adopts “context resets” to combat what we term Context Rot. The critical difference is that their Evaluator is itself an LLM (using Playwright browser automation for testing), whereas CAAF’s UAI is a deterministic assertion engine—a distinction that our Finding 9 demonstrates is decisive for constraint-satisfaction reliability: LLM-based evaluation achieves 0% paradox detection across 80 trials.
In the automotive safety domain, Zhang (2026) independently maps Harness Engineering onto the ISO 21448 (SOTIF) safety lifecycle, observing that feedforward constraints (ODD definitions) plus feedback monitoring (regression testing) in autonomous driving are structurally equivalent to the feedforward (Harness rules) plus feedback (UAI assertions) architecture of AI agent harnesses. This cross-domain convergence—AI engineering, industrial frontend development, and automotive safety engineering arriving at the same architectural pattern simultaneously—provides strong evidence that the Harness Engineering paradigm addresses a fundamental structural need rather than a domain-specific quirk.
@@ 12 Conclusion
[…]
@@ Appendix C Formal Deadlock Evidence Package
[SYSTEM DEADLOCK] Formal Paradox Report